OpenStudy (anonymous):

A particular strain of bacteria doubles in population every 10 minutes. Assuming you start with a single bacterium in a petri dish, how many bacteria will there be in 2.5 hours?

OpenStudy (anonymous):

So this equation is exponential... It will probably be in the form \[y=B ^{x}\]

OpenStudy (anonymous):

but y=B^x represents what

OpenStudy (anonymous):

Actually, the equation would be\[y = B ^ { \frac{ x }{ 10 }}\]

OpenStudy (anonymous):

B is the amount of bacteria made every ten minutes. X represents time.

OpenStudy (anonymous):

So b= 2

OpenStudy (anonymous):

Your equation is \[y = 2 ^ {\frac{ x }{ 10 }}\]

OpenStudy (anonymous):

great thank you

OpenStudy (anonymous):

Well I am not coming up with an answer that satisfies the choices I have

OpenStudy (anonymous):

So convert 2.5 hours to minutes

OpenStudy (anonymous):

That answer, in minutes, should be put in the X variable of the equation

OpenStudy (anonymous):

ahh thanks

OpenStudy (anonymous):

Ahh.....But still. Here. . 9536 B. 32,768 C. 65,536 D. 16,384 These are the choices. I do that equation and I got 30

OpenStudy (anonymous):

2.5 hours is 150 minutes 150/10 = 15 y = 2 ^ 15 y = 32,768

OpenStudy (anonymous):

It's 2^15, not 2x15

OpenStudy (anas.p):

The equation i gave you... \[a \times r^{n-1}\] Here r is 2 and n is 15, because 150 minutes divided by 10 and a is 1 so the equation becomes \[2^{150/10}\] which is exactly what @annikaroo said.

OpenStudy (anonymous):

Ahh I understand now thanks.